# A Tighter Complexity Analysis of SparseGPT

### v1

@@ A Tighter Complexity Analysis of SparseGPT
Xiaoyu Li xli216@stevens.edu. Stevens Institute of Technology.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song zsong@adobe.com, magic.linuxkde@gmail.com. Adobe Research.
In this work, we improved the analysis of the running time of SparseGPT [Frantar, Alistarh ICML 2023] from $O(d^{3})$ to $O(d^{\omega}+d^{2+a+o(1)}+d^{1+\omega(1,1,a)-a})$ for any $a\in[0,1]$, where $\omega$ is the exponent of matrix multiplication. In particular, for the current $\omega\approx 2.371$ [Alman, Duan, Williams, Xu, Xu, Zhou 2024], our running times boil down to $O(d^{2.53})$. This running time is due to the analysis of the lazy update behavior in iterative maintenance problems, such as [Deng, Song, Weinstein 2022, Brand, Song, Zhou ICML 2024].
@@ 1 Introduction
$O^{*}(d^{\omega}+d^{2+a}+d^{1+\omega(1,1,a)-a}).$
For the current $\omega\approx 2.371$, according to the Table 1 in [1], we should choose $a\approx 0.5275$ in order to balance the terms $d^{2+a+o(1)}$ and $d^{1+\omega(1,1,a)-a}$. Then the running time boils down to $O(d^{2.53})$, which is better than $O(d^{3})$ as claimed in [25].
@@ 2 Related Work
2.1 Model Acceleration and Pruning
Model acceleration is critical and urgently needed in many practical applications. There are many techniques for model acceleration. One line of work is to change model architecture to support fast inference, e.g., Mamba [27, 20], PolySketchFormer [48], Linearizing Transformers [86, 65], Hopfield Model [35, 34, 79, 83, 38, 80, 42] and so on. Another line of work is accelerating model computation on system level, e.g., FlashAttetnion [19, 18, 69, 74], parallel decoding [72], quantization [88, 78, 62] and so on. To accelerate LLMs training and inference, there is a line of work to approximate attention matrix computation [2, 3, 36, 87, 59, 67, 17, 58, 56, 29] in almost linear time. Some specific technique is developed to accelerate LLMs generation, e.g., KV-Cache compression [31, 51, 84, 90, 63, 22, 70] and speculative decodings [14, 50, 70, 23].
Pruning is a technique aimed at reducing the number of weights in a neural network by selectively removing certain neurons [52, 41]. This approach has gained considerable attention in recent years as a method to enhance the efficiency and scalability of deep learning models [40, 26, 49, 82, 5, 8, 64, 77, 13, 32, 33, 43, 24, 25, 71]. Pruning can be categorized based on its stage in the training process: pre-training pruning and post-training pruning. Pre-training pruning involves pruning the network at initialization. [26] demonstrated that a neural network pruned at initialization could be trained to achieve performance comparable to a dense model, a phenomenon referred to as the lottery ticket hypothesis. This discovery motivated a line of research focused on developing methods to reduce the computational cost of pruning neural networks at initialization [49, 82, 64, 77, 13]. More recently, [85] provided theoretical evidence that pre-training pruning can enhance a model’s generalization ability. Post-training pruning, initially popularized through its application in quantization [66, 39, 54, 71], was later extended to pruning by [33, 24, 47]. Post-training pruning aims to compress a well-optimized model using a small set of calibration data. This process typically involves layer-wise pruning of the neural network. Notably, [33] proposed a provable and efficient method for compressing a model by “stitching together” individually compressed layers.
@@ 2.2 Lazy Update
In recent years, the lazy update idea has emerged as a crucial technique to effectively and efficiently solve various optimization problems including linear programming [16, 11, 7, 46, 73, 61], semi-definite programming [44, 37, 28, 75], empirical risk minimization [60, 30, 68], cutting plane methods [45], neural network training [15, 10, 76], discrepancy minimization [21], dynamic attention problems [12] and so on.
We highlight several previous papers that share similar running time complexity. For comparison, we assume the input size of these problems is dominated by a parameter $d$. In the line of developments of fast linear program solvers, [16] first introduced the lazy update idea to develop efficient dynamic inverse structures for implementing interior point methods in solving linear programming problems, and it achieves the running time $O^{*}(d^{\omega}+d^{2.5-a/2}+d^{1.5+a})$. [73] uses a different method to achieve linear programming in the same complexity $O^{*}(d^{\omega}+d^{2.5-a/2}+d^{1.5+a})$ by incorporating the sparse sketching matrix to speed up the online matrix-vector multiplication. [46] improves the linear programming to $O^{*}(d^{\omega}+d^{2.5-a/2}+d^{1.5+a-\widetilde{a}/2}+d^{0.5+a+(\omega-1)\widetilde{a}})$ time where $\widetilde{a}\in[0,\alpha a]$ by designing a two-level lazy update framework for efficiently maintaining a projection matrix. Later [60] generalizes the algorithm for linear programming to empirical risk minimization, which also takes $O^{*}(d^{\omega}+d^{2.5-a/2}+d^{1.5+a})$ time. Different from [16], which employs a stochastic central path method that updates weights using a random sparse vector, [60] introduces a robust deterministic central path method. Additionally, [60] proposes an efficient data structure capable of handling updates efficiently even when the weight update vector is dense. Further, [68] uses a different method to achieve empirical risk minimization in the same complexity $O^{*}(d^{\omega}+d^{2.5-a/2}+d^{1.5+a})$ via online projection matrix-vector multiplication.
There are numerous other applications of the lazy update technique. For instance, [21] employs this approach to solve a subroutine of the discrepancy minimization problem in $O^{*}(d^{\omega}+d^{2+a}+d^{1+\omega(1,1,a)-a})$. This is achieved by designing a data structure that efficiently implements the iterative Edge-Walk partial-coloring algorithm proposed by [57], utilizing the lazy update concept. More recently, [12] demonstrates how the lazy update concept can be applied to a dynamic attention problem, achieving an amortized update time of $O(d^{\omega(1,1,a)-a})$.
@@ 3.1 Definitions and Facts
Algorithm 1 The SparseGPT algorithm (Algorithm 1 in [25]).
@@
3:     $\triangleright$ Weight matrix $W\in\mathbb{R}^{d\times d}$; Input feature matrix $X\in\mathbb{R}^{d\times d}$
4:     $\triangleright$ Lazy update block size $B\in\mathbb{N}_{+}$, $B=d^{a}$ for any $a\in[0,1]$; Adaptive mask size $B_{s}\in\mathbb{N}_{+}$
5:     $\triangleright$ Regularization parameter $\lambda>0$
6:     $M,~{}E\leftarrow{\bf 1}_{d\times d},~{}{\bf 0}_{d\times B}$triangleright $O(d^{2})$
7:     $\widetilde{H}\leftarrow(XX^{\top}+\lambda I_{d\times d})^{-1}$triangleright $O(d^{\omega})$ by Lemma 4.3
8:     for $i=0,B,2B,\ldots,\lfloor\frac{d}{B}\rfloor B$ do
9:         for $j=i+1,\ldots,i+B$ do
10:              if $j\bmod B_{s}=0$ then $\triangleright$ $O(d)$
11:                  $M_{*,[j,j+B_{s}]}\leftarrow\textsc{MaskSelect}(p,W_{*,[j,j+B_{s}]},\widetilde{H},j-1)$triangleright $O(d^{2}\log d)$ by Lemma 4.2
12:              end if
13:              $E_{*,j-i}\leftarrow(\mathbf{1}_{d\times 1}-M_{*,j})\circ W_{*,j}/\widetilde{H}_{j,j}$triangleright $O(d^{2})$
14:              $W_{*,[j,i+B]}\leftarrow W_{*,[j,i+B]}-E_{*,j-i}\widetilde{H}_{j,[j,i+B]}$triangleright $O(d^{2+a})$ by Lemma 4.4
15:         end for
16:         $W_{*,[i+B,d]}\leftarrow W_{*,[i+B,d]}-E\widetilde{H}_{[i,i+B],[i+B,d]}$triangleright $O(d^{1+\omega(1,1,a)-a})$ by Lemma 4.5
@@
18:     $W\leftarrow W\circ M$triangleright $O(d^{2})$
19:end procedure
20:
21:procedure MaskSelect($p\in[0,1]$, $W^{\prime}\in\mathbb{R}^{d\times r}$, $\widetilde{H}\in\mathbb{R}^{d\times d},s\in\mathbb{N}_{+}$)
22:     $\triangleright$ Sub-weight matrix $W^{\prime}\in\mathbb{R}^{d\times r}$; Inverse of Hessian matrix $\widetilde{H}\in\mathbb{R}^{d\times d}$
23:     $\triangleright$ Index $s\in\mathbb{N}_{+}$, recording the position of $W^{\prime}$ in $W$
24:     $M^{\prime}\leftarrow\mathbf{0}_{d\times r}$triangleright $O(dr)$
25:     for $k=1,\ldots,r$ do
26:         $w\leftarrow W^{\prime}_{*,k}$triangleright $w\in\mathbb{R}^{d}$, $O(dr)$
27:         $w\leftarrow(w\circ w)/(\widetilde{H}_{s+k,s+k})^{2}$triangleright $O(dr)$
28:         $J\leftarrow$ indices of top $(1-p)d$ largest entries of $w$triangleright $O(r\cdot d\log d)$ by sorting
29:         for $j\in J$ do
30:              $M^{\prime}_{k,j}\leftarrow 1$triangleright $O(dr)$
31:         end for
32:     end for
33:     return $M^{\prime}$
34:end procedure
@@ 4 Main Results
  • Line 6 takes time $O(d^{2})$ to initiate $M$ and takes time $O(d^{1+a})$ to initiate $E$.
  • By Lemma 4.3, Line 7 takes time $O(d^{\omega})$.
  • In each iteration, Line 10 takes time $O(1)$ to check if $j\bmod B_{s}=0$. Since there are $d$ iterations, the total time is $O(d)$.
  • By Lemma 4.2, Line 11 takes time $O(d^{2}\log d)=O(d^{2+o(1)})$ over all iterations.
  • In each iteration, Line 13 takes time $O(d)$ to compute $W_{*,j}/[\widetilde{H}]_{jj}^{2}$ and takes time $O(d)$ to compute $1_{d\times 1}-M_{*,j}$ and then takes time $O(d)$ to compute the Hadamard product of them. Since there are $d$ iterations, the total time is $O(d^{2})$.
  • By Lemma 4.4, Line 14 takes time $O(d^{2+a})$ over all iterations.
  • By Lemma 4.5, Line 16 takes time $O(d^{1+\omega(1,1,a)-a})$ over all iterations.
  • Line 18 takes time $O(d^{2})$ to compute the Hadamard product of two $d\times d$ matrices.
@@
4.1 Full Complexity Analysis
@@
The running time of Procedure MaskSelect in Algorithm 1 is $O(rd\log d)$. Hence the running time of Line 11 in Algorithm 1 over all iterations is $O(d^{2}\log d)$.
@@
  • Line 24 takes time $O(dr)$ to initialize $M^{\prime}$.
  • In every iteration, Line 26 takes time $O(d)$ to initialize $w$. Since there are $r$ iterations, the total running time is $O(dr)$.
  • In every iteration, Line 27 takes time $O(1)$ to compute $(\widetilde{H}_{s+k})^{2}$ and then takes time $O(d)$ to compute $(w\circ w)/a$. Since there are $r$ iterations, the total running time is $O(dr)$.
  • In every iteration, Line 28 takes time $O(d\log d)$ to sort $w$ (without overwriting) and takes time $O(d)$ to read the indices of top $(1-p)d$ largest entries. Since there are $r$ iterations, the total running time is $O(r(d\log d+d))=O(rd\log d)$.
  • In every iteration over $k$, 30 takes time $O(d)$ to update $M^{\prime}$. Since there are $r$ iterations, the total running time is $O(dr)$.
@@
Now, we analyze the running of Line 11 over all iterations in Algorithm 1. Let $r=B_{s}$. The total number calls to MaskSelect is $d/B_{s}$. Hence the running over all iterations is
@@
The running time of Line 7 in Algorithm 1 is $O(d^{\omega})$.
@@
Line 7 takes time $O(d^{\omega})$ to compute $XX^{\top}$, takes time $O(d)$ to add $XX^{\top}$ and $\lambda I_{d\times d}$. Then computing the inverse of a $d\times d$ matrix takes time $O(d^{\omega})$. Hence Line 7 takes time
@@
The running time of Line 14 in Algorithm 1 over all iterations is $O(d^{2+a})$.
@@
In every iteration, Line 14 involves one matrix addition and one matrix multiplication. The size of $E_{*,j-i}$ is $d\times 1$ and the size of $\widetilde{H}_{j,[j,i+B]}$ is $1\times B$. Hence multiplying $E_{*,j-i}$ with $\widetilde{H}_{j,[j,i+B]}$ takes time $O(dB)$. Subtracting $E_{*,j-i}\widetilde{H}_{j,[j,i+B]}$ from $W_{*,[j,i+B]}$ takes time $O(dB)$. Since there are $d$ iterations, the running time of Line 14 in Algorithm 1 over all iterations is
@@
The running time of Line 16 in Algorithm 1 over all iterations is $O(d^{1+\omega(1,1,a)-a})$.
@@
In every iteration, Line 16 involves one matrix addition and one matrix multiplication. The size of $E$ is $d\times B$. The size of $\widetilde{H}_{[i,i+B],[i+B,d]}$ is $B\times(d-(i+B))$. Multiplying $E$ with $\widetilde{H}_{[i,i+B],[i+B,d]}$ takes time ${\cal T}_{\mathrm{mat}}(d,B,d-(i+B))={\cal T}_{\mathrm{mat}}(d,B,d)$. Subtracting $E\widetilde{H}_{[i,i+B],[i+B,d]}$ from $W_{*,[i+B,d]}$ takes time $O(dB)$. Hence, the running time of Line 16 over all iterations is

### v2

@@ A Tighter Complexity Analysis of SparseGPT
Xiaoyu Li xli216@stevens.edu. Stevens Institute of Technology.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.
In this work, we improved the analysis of the running time of SparseGPT [Frantar, Alistarh ICML 2023] from $O(d^{3})$ to $O(d^{\omega}+d^{2+a+o(1)}+d^{1+\omega(1,1,a)-a})$ for any $a\in[0,1]$, where $\omega$ is the exponent of matrix multiplication. In particular, for the current $\omega\approx 2.371$ [Alman, Duan, Williams, Xu, Xu, Zhou 2024], our running time boils down to $O(d^{2.53})$. This running time is due to the analysis of the lazy update behavior in iterative maintenance problems such as [Deng, Song, Weinstein 2022; Brand, Song, Zhou ICML 2024].
@@ 1 Introduction
$O(d^{\omega}+d^{2+a+o(1)}+d^{1+\omega(1,1,a)-a}).$
For the current $\omega\approx 2.371$, according to the Table 1 in ADW+ [2], we should choose $a\approx 0.5275$ in order to balance the terms $d^{2+a+o(1)}$ and $d^{1+\omega(1,1,a)-a}$. Then the running time boils down to $O(d^{2.53})$, which is better than $O(d^{3})$ as claimed in FA [29].
@@ 2 Related Work
2.1 Model Acceleration
Model acceleration is critical and urgently needed in many practical applications. There are many techniques for model acceleration. One line of work is to change model architecture to support fast inference, e.g., Mamba [31, 23], PolySketchFormer [54], Linearizing Transformers [99, 97, 74], Hopfield Model [39, 38, 89, 93, 42, 90, 48, 46] and so on. Another line of work is accelerating model computation on system level, e.g., FlashAttetnion [22, 21, 78, 84, 62], parallel decoding [82], quantization [101, 88, 71, 98] and so on. To accelerate LLMs training and inference, there is a line of work to approximate attention matrix computation [3, 4, 40, 100, 68, 76, 20, 67, 63, 33, 67, 66, 16, 47, 19] in almost linear time. Some specific technique is developed to accelerate LLMs generation, e.g., KV-Cache compression [35, 57, 94, 103, 72, 26, 79, 81] and speculative decodings [15, 56, 79, 27, 24].
2.2 Model Pruning
Pruning is a technique aimed at reducing the number of weights in a neural network by selectively removing certain neurons [58, 45]. This approach has gained considerable attention in recent years as a method to enhance the efficiency and scalability of deep learning models [44, 30, 55, 92, 6, 9, 73, 87, 14, 36, 37, 49, 28, 29, 80]. Pruning can be categorized based on its stage in the training process: pre-training pruning and post-training pruning. Pre-training pruning involves pruning the network at initialization. FC [30] demonstrated that a neural network pruned at initialization could be trained to achieve performance comparable to a dense model, a phenomenon referred to as the lottery ticket hypothesis. This discovery motivated a line of research focused on developing methods to reduce the computational cost of pruning neural networks at initialization [55, 92, 73, 87, 14]. More recently, YLG+ [96] provided theoretical evidence that pre-training pruning can enhance a model’s generalization ability. Post-training pruning, initially popularized through its application in quantization NAVB+ [75], HNH+ [43], LGT+ [60], SLBK [80], was later extended to pruning by HCI+ [37], FA [28], KKM+ [53], LLS+24c [64]. Post-training pruning aims to compress a well-optimized model using a small set of calibration data. This process typically involves layer-wise pruning of the neural network. Notably, HCI+ [37] proposed a provable and efficient method for compressing a model by “stitching together” individually compressed layers.
@@ 2.3 Lazy Update
In recent years, the lazy update idea used in iterative maintenance problems has emerged as a crucial technique to effectively and efficiently solve various optimization problems, including linear programming [18, 12, 8, 52, 83, 70], semi-definite programming [50, 41, 32, 85], empirical risk minimization [69, 34, 77], cutting plane methods [51], neural network training [17, 11, 86], discrepancy minimization [25], dynamic attention problems [13] and so on.
We highlight several previous papers that share similar running time complexity. For comparison, we assume the input size of these problems is dominated by a parameter $d$. In the line of developments of fast linear program solvers, CLS [18] first introduced the lazy update idea to develop efficient dynamic inverse structures for implementing interior point methods in solving linear programming problems, and it achieves the running time $O^{*}(d^{\omega}+d^{2.5-a/2}+d^{1.5+a})$. SY [83] uses a different method to achieve linear programming in the same complexity $O^{*}(d^{\omega}+d^{2.5-a/2}+d^{1.5+a})$ by incorporating the sparse sketching matrix to speed up the online matrix-vector multiplication. JSWZ [52] improves the linear programming to $O^{*}(d^{\omega}+d^{2.5-a/2}+d^{1.5+a-\widetilde{a}/2}+d^{0.5+a+(\omega-1)\widetilde{a}})$ time where $\widetilde{a}\in[0,\alpha a]$ by designing a two-level lazy update framework for efficiently maintaining a projection matrix. Later, LSZ [69] generalizes the algorithm for linear programming to empirical risk minimization, which also takes $O^{*}(d^{\omega}+d^{2.5-a/2}+d^{1.5+a})$ time. Different from CLS [18], which employs a stochastic central path method that updates weights using a random sparse vector, LSZ [69] introduces a robust deterministic central path method. Additionally, LSZ [69] proposes an efficient data structure capable of handling updates efficiently even when the weight update vector is dense. Further, QSZZ [77] uses a different method to achieve empirical risk minimization in the same complexity $O^{*}(d^{\omega}+d^{2.5-a/2}+d^{1.5+a})$ via online projection matrix-vector multiplication.
There are numerous other applications of the lazy update technique. For instance, DSW [25] employs this approach to solve a subroutine of the discrepancy minimization problem in $O^{*}(d^{\omega}+d^{2+a}+d^{1+\omega(1,1,a)-a})$. This is achieved by designing a data structure that efficiently implements the iterative Edge-Walk partial-coloring algorithm proposed by LM [65], utilizing the lazy update concept. More recently, BSZ [13] demonstrates how the lazy update concept can be applied to a dynamic attention problem, achieving an amortized update time of $O(d^{\omega(1,1,a)-a})$.
@@ 3.2 Definitions and Facts
Algorithm 1 The SparseGPT algorithm (Algorithm 1 in FA [29]).
@@
3:     $\triangleright$ Weight matrix $W\in\mathbb{R}^{d\times d}$
4:     $\triangleright$ Input feature matrix $X\in\mathbb{R}^{d\times d}$
5:     $\triangleright$ Lazy update block size $B\in\mathbb{N}_{+}$, $B=d^{a}$ for any $a\in[0,1]$
6:     $\triangleright$ Adaptive mask size $B_{s}\in\mathbb{N}_{+}$
7:     $\triangleright$ Regularization parameter $\lambda>0$
8:     $M,~{}E\leftarrow{\bf 1}_{d\times d},~{}{\bf 0}_{d\times B}$triangleright $O(d^{2})$
9:     $\widetilde{H}\leftarrow(XX^{\top}+\lambda I_{d\times d})^{-1}$triangleright $O(d^{\omega})$ by Lemma 5.2
10:     for $i=0,B,2B,\ldots,\lfloor\frac{d}{B}\rfloor B$ do
11:         for $j=i+1,\ldots,i+B$ do
12:              if $j\bmod B_{s}=0$ then $\triangleright$ $O(d)$
13:                  $M_{*,[j,j+B_{s}]}\leftarrow\textsc{MaskSelect}(p,W_{*,[j,j+B_{s}]},\widetilde{H},j-1)$triangleright $O(d^{2}\log d)$ by Lemma 5.1
14:              end if
15:              $E_{*,j-i}\leftarrow(\mathbf{1}_{d\times 1}-M_{*,j})\circ W_{*,j}/\widetilde{H}_{j,j}$triangleright $O(d^{2})$
16:              $W_{*,[j,i+B]}\leftarrow W_{*,[j,i+B]}-E_{*,j-i}\widetilde{H}_{j,[j,i+B]}$triangleright $O(d^{2+a})$ by Lemma 5.3
@@
18:         $W_{*,[i+B,d]}\leftarrow W_{*,[i+B,d]}-E\widetilde{H}_{[i,i+B],[i+B,d]}$triangleright $O(d^{1+\omega(1,1,a)-a})$ by Lemma 5.4
19:     end for
20:     $W\leftarrow W\circ M$triangleright $O(d^{2})$
21:end procedure
22:
23:procedure MaskSelect($p\in[0,1]$, $W^{\prime}\in\mathbb{R}^{d\times r}$, $\widetilde{H}\in\mathbb{R}^{d\times d},s\in\mathbb{N}_{+}$)
24:     $\triangleright$ Sub-weight matrix $W^{\prime}\in\mathbb{R}^{d\times r}$; Inverse of Hessian matrix $\widetilde{H}\in\mathbb{R}^{d\times d}$
25:     $\triangleright$ Index $s\in\mathbb{N}_{+}$, recording the position of $W^{\prime}$ in $W$
26:     $M^{\prime}\leftarrow\mathbf{0}_{d\times r}$triangleright $O(dr)$
27:     for $k=1,\ldots,r$ do
28:         $w\leftarrow W^{\prime}_{*,k}$triangleright $w\in\mathbb{R}^{d}$, $O(dr)$
29:         $w\leftarrow(w\circ w)/(\widetilde{H}_{s+k,s+k})^{2}$triangleright $O(dr)$
30:         $J\leftarrow$ indices of top $(1-p)d$ largest entries of $w$triangleright $O(r\cdot d\log d)$ by sorting
31:         for $j\in J$ do
32:              $M^{\prime}_{k,j}\leftarrow 1$triangleright $O(dr)$
33:         end for
34:     end for
35:     return $M^{\prime}$
36:end procedure
@@ 4 Main Results
  • Line 8 takes time $O(d^{2})$ to initiate $M$ and takes time $O(d^{1+a})$ to initiate $E$.
  • By Lemma 5.2, Line 9 takes time $O(d^{\omega})$.
  • In each iteration, Line 12 takes time $O(1)$ to check if $j\bmod B_{s}=0$. Since there are $d$ iterations, the total time is $O(d)$.
  • By Lemma 5.1, Line 13 takes time $O(d^{2}\log d)=O(d^{2+o(1)})$ over all iterations.
  • In each iteration, Line 15 takes time $O(d)$ to compute $W_{*,j}/[\widetilde{H}]_{jj}^{2}$ and takes time $O(d)$ to compute $1_{d\times 1}-M_{*,j}$ and then takes time $O(d)$ to compute the Hadamard product of them. Since there are $d$ iterations, the total time is $O(d^{2})$.
  • By Lemma 5.3, Line 16 takes time $O(d^{2+a})$ over all iterations.
  • By Lemma 5.4, Line 18 takes time $O(d^{1+\omega(1,1,a)-a})$ over all iterations.
  • Line 20 takes time $O(d^{2})$ to compute the Hadamard product of two $d\times d$ matrices.
@@
Remark 4.2.
The $O(d^{2})$ space complexity is sufficient to implement fast matrix multiplication, which is equivalent to that of standard matrix multiplication. In fact, all algorithms derived from Strassen’s original approach exhibit a $\Theta(d^{2})$ space complexity. For further details, we refer readers to [95, 1].
5 Detailed Complexity Analysis
@@
The running time of Procedure MaskSelect in Algorithm 1 is $O(rd\log d)$. Hence the running time of Line 13 in Algorithm 1 over all iterations is $O(d^{2}\log d)$.
@@
  • Line 26 takes time $O(dr)$ to initialize $M^{\prime}$.
  • In every iteration, Line 28 takes time $O(d)$ to initialize $w$. Since there are $r$ iterations, the total running time is $O(dr)$.
  • In every iteration, Line 29 takes time $O(1)$ to compute $(\widetilde{H}_{s+k})^{2}$ and then takes time $O(d)$ to compute $(w\circ w)/a$. Since there are $r$ iterations, the total running time is $O(dr)$.
  • In every iteration, Line 30 takes time $O(d\log d)$ to sort $w$ (without overwriting) and takes time $O(d)$ to read the indices of top $(1-p)d$ largest entries. Since there are $r$ iterations, the total running time is $O(r(d\log d+d))=O(rd\log d)$.
  • In every iteration over $k$, Line 32 takes time $O(d)$ to update $M^{\prime}$. Since there are $r$ iterations, the total running time is $O(dr)$.
@@
Now, we analyze the running of Line 13 over all iterations in Algorithm 1. Let $r=B_{s}$. The total number calls to MaskSelect is $d/B_{s}$. Hence the running over all iterations is
@@
The running time of Line 9 in Algorithm 1 is $O(d^{\omega})$.
@@
Line 9 takes time $O(d^{\omega})$ to compute $XX^{\top}$, takes time $O(d)$ to add $XX^{\top}$ and $\lambda I_{d\times d}$. Then computing the inverse of a $d\times d$ matrix takes time $O(d^{\omega})$. Hence Line 9 takes time
@@
The running time of Line 16 in Algorithm 1 over all iterations is $O(d^{2+a})$.
@@
In every iteration, Line 16 involves one matrix addition and one matrix multiplication. The size of $E_{*,j-i}$ is $d\times 1$ and the size of $\widetilde{H}_{j,[j,i+B]}$ is $1\times B$. Hence multiplying $E_{*,j-i}$ with $\widetilde{H}_{j,[j,i+B]}$ takes time $O(dB)$. Subtracting $E_{*,j-i}\widetilde{H}_{j,[j,i+B]}$ from $W_{*,[j,i+B]}$ takes time $O(dB)$. Since there are $d$ iterations, the running time of Line 16 in Algorithm 1 over all iterations is
@@
The running time of Line 18 in Algorithm 1 over all iterations is $O(d^{1+\omega(1,1,a)-a})$.
@@
In every iteration, Line 18 involves one matrix addition and one matrix multiplication. The size of $E$ is $d\times B$. The size of $\widetilde{H}_{[i,i+B],[i+B,d]}$ is $B\times(d-(i+B))$. Multiplying $E$ with $\widetilde{H}_{[i,i+B],[i+B,d]}$ takes time ${\cal T}_{\mathrm{mat}}(d,B,d-(i+B))={\cal T}_{\mathrm{mat}}(d,B,d)$. Subtracting $E\widetilde{H}_{[i,i+B],[i+B,d]}$ from $W_{*,[i+B,d]}$ takes time $O(dB)$. Hence, the running time of Line 18 over all iterations is